Scraps of text for incorporation in the VIS Meeting Minutes

KAGRA/Subgroups/VIS/MeetingNAOJ

Type-B

Type-Bp

PRM

According to Tomaru-san's measurements of the chamber height with respect to the BS, it is necessary to lower the suspension by 6.3 mm. We checked whether this is possible using the SF keystone:

PR3

Hirata-san finished the design of the SF keystone ballast mass base plate. We will proceed as soon as we have the new components.

Type-B and Type-Bp height report

Type-B

Type-Bp

Type-B and Type-Bp various

Type-Bp PR3

Type-B

Type-Bp PR3

Type-B

Type-Bp

SR3

Type-B

Type-Bp

On Monday we double-checked the problems that were detected during the cable health check (klog 18240): we need to check PR3 BF LVDTs H3 and V3, which have higher resistances than expected in one of their coils.

Type-Bp

Type-B

With the help of lots of people we finished checking the cable health in Type-B suspensions. The highlights are:

Type-Bp

We assessed the health of the in-vacuum cables by measuring resistances, capacitances and inductances from the outside of the chamber at the flanges.

Type-B

[Fabián]

[Terrence]

  1. Fabian set up the optical lever for SR3 and calibrated the QPDs https://klog.icrr.u-tokyo.ac.jp/osl/?r=17875.

  2. I implemented the calibration factors and I derived the sensing matrix OPLEV2EUL using old data and diagonalized the optical levers using the SENSALIGN matrix. https://klog.icrr.u-tokyo.ac.jp/osl/?r=17914.

  3. I reset the sensing matrix LVDT2EUL and actuation matrix COIL2EUL at the IP using Marks's old derivation because I changed it two years ago and I don't think they are reliable. Note that I didn't do diagonalization for these. https://klog.icrr.u-tokyo.ac.jp/osl/?r=17868.

  4. I measured and fit transfer functions for all 15 degrees of freedom. (klog pending). All models were updated in the control filter banks as FM10.
  5. I made control filters base on some simple criteria (IP: DCCTRL, GAS: DAMP, IM: DAMP, TM: OLDAMP). (klog pending)
    • Filter bank convention: suspension model in FM10, -1 in FM9, DAMP filter in FM4, and if applicable, DC filters in FM3, and notches at FM5.
    • We don't have parallel control banks to separate DC and damping control. These DC filters work together with the DAMP filter and are not expected to work alone although it's stable to the DC filters turned out without the DAMP filters. DAMP filters, on the other hand, can work alone for damping.
  6. As of now, all loops are closed for 4 hours+ and work fine without issues. Furthermore, it survived the Vanuatu 6.8 magnitude earthquake a few hours ago (all seismic monitors went red at KAGRA).

Something to note 1. LVDTs of the GAS filters are very noisy especially at above 10 Hz. I see strange responses above 10 Hz. I have made additional low-pass filters for these controls.

  1. There's no windshield for the optical lever. So, the optical lever readouts, especially longitudinal, can be very noisy.
  2. I observed that the yaw chain mode was not very well damped. This mode is supposed to be damped by the passive damper. We can see very slow yaw oscillation right after the IP yaw controls are engaged.
  3. There's no inertial damping. I didn't touch the IDAMP filter bank and the BLEND filter bank at the IP.
  4. There's no IM2TM DC control. I can implement this if this is requested. However, note that there's no dedicated setpoint for this. Alternatively, the OPTICALIGN block can be used to offset the IM and hence the TM for alignment.
  5. Guardian has changed a lot and it's not maintained by me anymore. I don't how what's the current situation so I did not test it.
  6. Everything is reproducible from my jupyter notebook at /kagra/Dropbox/Subsystems/VIS/vis_commissioning/sr3/notebook.

Type-B

Type-Bp

Type-B

Type-Bp

PR3:

Type-B

SR3

BS

For the O3GK paper, I revisited the calculation of the effect of DAC and coil driver noise on DARM. Because no one measured the transfer function from individual coils to DARM we cannot propagate the noise directly. However, it is possibler to calculate upper limits using the transfer functions from L, P and Y of the mirrors.

Type-Bp

Type-B

SRM

BS height

BS supporting beams

Type-B

I wrote the Type-B section of the O3GK paper.

BS

SR2

We measured the IP transfer functions. The resonance frequencies in L and T are both close to 70 mHz. They cannot be distinguished from each other but they below 80 mHz.

SR3

Remaining tasks

Type-Bp

Type-B

Limit switches in SR2 and SR3

There were three switches that didn't seem to be working. Hirata-san fixed them and we checked them with the multimeter. They are working now

SRM

BS

We revisited the BS IP to check stability and values of resonance frequencies.

Type-B

Activity report

What's left to do?

We aim to finishe next week

SR2

SR3

Type-B BS

Type-Bp

Type-Bp

PR2

PRM

Type-Bp

PR2

As people may remember we had to move the SD keystone very high in order to have the mirror at the appropriate height:

PR3

Yet to do in PRM

PR2

On Monday:

On Tuesday

On Wednesday

PRM

Type-Bp PR3

Type-Bp remaining work

Type-Bp general

Type-Bp PR3

Type-Bp PR2

Type-Bp PR2

This week we inspected and tightened the wires clapms in PR2. We followed the same procedure as in PR3, so I won't write any details here.

Type-Bp PR3

This week we did two tasks, both related with PR3 clamps:

Size of gaps

Gluing test.

Is there another option before deciding to hang the mirror again? Yes

Type-Bp PR3

Given the fact that the gluing of the wire underneath the mirror didn't fix the problem of the stability of the mirror, we decided to tighten the wire clamps.

Type-Bp PR3

Klog

Signal generator recovery

I set the amplitude and frequencies of the signal generators according to the values written on the devices themselves or the klog entries:

PR3

In ordert to check the health of the IM-Y picomotor we released the IM from touching the IRM temporarily as follows:

This suggested we could use the picomotor to move the IM from around -3 mrad back to zero.

We modified the clamping of a cable that was almost touching another cable.

X End

Type-Bp

Type-Bp

Klog draft

With Hirata-san, Washimi-san and Ikeda-san.

Another klog draft

I moved the BF to a position similar to the one in which the BF was before the BF picomotor was moved originally. The idea is to check whether the alignment of the IM is recovered.

Result: the alignment of the IM was not recovered.

Before moving the BF picomotors the pisition of the BF was

I used the BF picomotor 2 and some BF Damper coil magnet actuator to move it to

The position is not exaclty the same but it's very similar. The IM alignment did not improve, the IM-Y magnitude is still large:

See this notebook page for more details.

Type-B

Albums

SR3 ballast mass adjustment highlights

On top of Intermediate Mass

On top of Bottom Filter (klog yet to be written)

A klog: Ballast mass adjustment on top of BF.

With Hirata-san and Washimi-san

SR3 remedying work after O3

We need to adjust the amount of ballast mass on tip of BF because F1 fishing rod is jammed. Today we tested how mass we need to add.

Type-B

Albums

SR3 alignment recovery (klog15928)

Plan for next week

Type-Bp

Albums

PR3

Gluing of the mirror wire

Traverser in PR3 and PR2

-----------------------

General

[Fabian, Hirata] We inspected SR3 payload and chamber and came up with a plan for installing a thermistor.

Type-B

Albums

Type-Bp

Albums

PR3

Type-B

Albums

Week of the 18th of January

SRM

I disconnected SRM picomotor drivers to move them away from the assembly frame. I failed to report it in a klogin due time.

BS

Plan for next week

SR3

Type-B

Albums

Week of the 11th of January

BS: klog 15729

SR3

We moved the SR3 small rack inside of the clean booth to a new permanent position.See this the location in this picture and this other one.

Plan for next week

Type-B

Albums

Week of the 4th of January

SR3

BS

Type-B all

We checked the flanges we could use for the cables of a thermistor. See details in klog 15700.

Type-B

Albums

Week of the 14th of December

Week of the 21st of December

Type-B

Albums

Week of the 7th of December

Type-B

Week of the 30th of November

General

BS

SRM, SR3 and SR2

A TO DO list for SR suspensions.

Type-B

Week of the 23rd of November 2020

[Hirata, Washimi, Satou, Aso, Fabian]

SRM

SR3

Type-B

[Hirata-san, Washmi-san and Satou-san, Ikeda-san, Fabian]

General

SR3

SRM trouble mechanical troubleshooting

Netx two weeks plan

Notes for SRM Klog on 19-11-2020

Assessment of BF and IM

Conclusion

Conditions

Description of experiments

  1. The BF keystone is at 288 um and the IM-V at 321 um.
    • The BF transfer function looks very good (picture 1p1).
    • The IM-V transfer function looks also good (picture 1p2). The vertical OSEMs don't seem to be out of range although they might be close to their limit. I need to check.
    • The IM-Y transfer function also looks good (picture 1p3).
    • When it swings free there is no large oscillation (picture 1p4).
  2. BF keystone down to 172 um (-2000 cnt of actuation).
    • A large oscillation appears: 8 um-pp at 30 mHz. See 2p2.
    • The BF transfer function looks very good (picture 2p3).The resonant frequency is at 273 mHz.
    • The IM-V transfer function looks also good (picture 2p4).
  3. I moved the BF keystone down to 132 um with -2800 cnt of actuation. IM-V goes down to 185 um.
    • BF oscillates with an amplitude of 9 um-pp at about 30 mHz, but IM-V does not move! See picture 3p1.
    • Then I applied a sinusoidal actuation with an amplitude of 2000 counts at 100 mHz. See picture 3p2, 3p3. I removed the AC actuation and measured transfer functions.
    • IM-Y and IM-L transfer functions looks fine. See 3p4, 3p5.
    • IM-P looks good. See 3p6.
    • IM-R looks good although noisy at high frequencies. See 3p7.
    • IM-V looks really bad. See 3p8.
  4. Then I set the DC actuation to -2000 counts and turned the AC actuation on. The IM vertical OSEMs seem to be completely free from any stop and follow the actuation. Pictures 4p1, 4p2, 4p3. The BF follows the actuation also.
  5. I removed the sinusoidal actuation completely and kept the AC one. The BF oscillates with a very low frequency again. See 5p1, 5p2.
  6. I removed the DC actuation completely. The low frequency oscillation in BF is gone. See 6p1.

Assessment of F1

Conclusions

It seems that when the keystone is placed around zero (original nominal) and above, a coupling with other mechanical elements appear. This manifests itself as persistent oscillation above 310 mHz, the decrease in the height of a peak in transfer functions, change in its frequency and a decrease of coherence below such a frequency. The system seems to be free of stray couplings at -572 um or below. However, a clear boundary between healthy and unhealthy behaviour has not been established yet.

Conditions

Desription of measurements

  1. I left F1 keystone at -182 um (10,000 counts of actuation) free swinging last night and, despite the time elapsed, the oscillation at 312.5 mHz has not damped (picture 1p1, 1p2). It has an amplitude of 30 um-pp. I assume it continues because the excitation hasn't ceased. The transfer function looks very good (1p3) but the resonant frequency is at 273 mHz not at 312.5 mHz.

  2. I removed the actuation and the keystone goes down to -572 um. The large oscillation at 312.5 mHz is gone; we see the highest peak at the resonant frequency of 281 mHz (pictures 2p1, 2p2, 1p3). The transfer function looks very good. The peak shows up at 273 mHz not at 281 mHz (picture 2p4).

  3. I turned the actuation on again (10,000 counts, keystone at around -182 um). The large oscillation is back, this time at 304 mHz, together with other oscillations in that frequency neighborhood (pictures 3p1 and 3p2). Then I turned the control system on, damping only, but it didn't damp it. I increased the gain but it became unstable. I don't have the TF in Foton to check the UGF and gain, so I gave up damping it.
  4. I increased the actuation to 14,400 counts to bring the keystone around zero. The oscillation is still there (pictures 4p1, 4p2) and it appears in a slightly higher frequency, at 328 mHz. The transfer function shows an unhealthy system: the peak around 280 mHz is gone and a shorter peak appears at 304 mHz instead (picture 4p3). The coherence at that frequency is also high for a resonance (0.91).The coherence at low frequencies does not look as good.

  5. I increased the actuation to 25,000 counts, which brought the keystone to around 450 um. The large oscillation is there (pictures 4p1, 4p2) at 320 mHz. The transfer function shows again an unhealthy system with the resonance at 320 mHz and with lower coherence below the first resonance (picture 5p3).

Klog for 16-11-2020

With Hiarata-san and Washimi-san.

Preparation from the control room

In situ preparation

Work inside the chamber

Type-B

With Washimi-san.

Type-Bp

With Washimi-san

General

About LVDT cards

[Fabian and Washim-san]

With the new information that Alessadro (Nikhef) provided a few weeks ago, we began measuring the settings in PR suspensions.

Type-B

[Fabian and Washimi-san]

Type-Bp

[Fabian and Washimi-san]

Type-B

[Fabian]

Possible activities for the next weeks (subject to Takahashi-san's approval):

Type B

Report for the week of the 19th of October 2020.

[Fabian]

General

Type B

Report of the week of the 12th of October

Type B

Report of the week of the 28th of Septermber

Type B

Report for the week of the 7th of September 2020.

Type B

Report for the week of the 27th of April 2020.

In the context of the Type-B paper I checked the noise level of the gephones: I did a calculation of the expected sensitivity and campared with Fujii-kun's measurements:

In the context of the writing

Type B

Report

Report

Report for the week 16-12-2019

Plan for next week

Report for the week 09-12-2019

Plan for next week

Vacation/Travel Info

Type B

Report for the week 23-09-2019

Plan for next week

Vacation/Travel Info

Report for the week 23-09-2019

Plan for next week

Vacation/Travel Info

Type B

Report for the week 13-09-2019

Type B

Report for the week 02-09-2019

Plan for next week

Report for the week 26-08-2019 (Written by Fabian)

Plan for the week 02-09-2019

Type B

Report for the week 15-08-2019 (Written by Fabian)

Plan for the week 19-08-2019

Type B

Report for the week 05-08-2019 (Written by Fabian)

Plan for the week 15-08-2019

Report for the week 7/29 (Written by Fabian)

Plan for the week 8/5

Type B

Report for the week 7/22 (Written by Fabian)

Plan for the week 7/29

Vacation/Travel Info

Type B (Mark)

Report for the week 7/15 (Written by Fabian)

Plan for the week 7/22

Vacation/Travel Info

Type B (Mark)

Report for the week 7/8 (Written by Fabian)

Plan for the week 15/8

Vacation/Travel Info

Type B (Mark)

Report for the week 7/1 (Written by Fabian)

Plan for the week 7/8

Type B (Mark)

Report for the week 6/24 (Written by Terrence and Fabian)

Plan for the week 7/1

Vacation/Travel Info

Type B (Mark)

Report for the week 6/17

Plan for the week 6/24

Vacation/Travel Info

Type B (Mark)

Report for the week 6/10

BS

Type B (Mark)

Report for the week 6/3

SRM Inertial Damping and Residual Motion Stuff

RMS Displacement:

Requirement

Result

L

0.4 µm

0.1076 µm

P

1 µrad

0.0569 µrad

Y

1 µrad

0.0598 µrad

RMS Velocity:

Requirement

Result

L

0.5 µm

0.1014 µm/s

P

/

0.2165 µrad/s

Y

/

0.1776 µrad/s

SR3 Inertial Damping

Type B (Mark)

Report for the week 5/27 (written by Fabian)

Type B (Mark)

Report for the week 5/20 (written by Fabian)

Plan for the week 5/25

Report for Golden Week

Report for the week of 5/6

Report for week of 5/13

Plan for week of 5/20

Vacation/Travel Info

KAGRA/Subgroups/VIS/TypeB/Minutes (last edited 2021-10-22 10:29:26 by fabian.arellano)